# A Real-time Dyson Expansion Scheme: Efficient Inclusion of Dynamical Correlations in Non-equilibrium Spectral Properties

### v1

@@ Introduction
In Fig. 3 b) we again show the time resolved emission spectrum, now computed using the HF-GKBA. This result shares features common to TD-HF as we expect because it is dominated by the off-diagonal terms dominated by HF. Firstly, while the HF signatures are diminished, we see a similar stretching of the bands relative to their equilibrium value. One reason for this change comes from an inconsistency between the treatment of the time diagonal and off-diagonal terms in HF-GKBA. Specifically it arises from the GKBA density matrix, $\rho^{\mathrm{GKBA}}(t)$, entering the HF equation for the off-diagonal instead of $\rho^{\mathrm{HF}}(t)$. When solving the TD-HF equation the density matrix will be projected on to a set of excited states of the HF Hamiltonian. Since in general at time $t$, $\rho^{\mathrm{GKBA}}(t)\neq\rho^{\mathrm{HF}}(t)$ the set of excited states we project onto will be different thus leading to a distinct energy spectrum. Such an inconsistency makes it a priori difficult to interpret HF-GKBA spectral results or to say how the difference from TD-HF will manifest. Another important observation is the lack of additional excitonic features. The picture is thus consistent with TD-HF (albeit with the magnitude of the dynamical band renormalization being significantly distinct). This is expected as the spectrum is dominated by the time off-diagonal terms treated at the HF level.
@@
Finally we turn to Fig. 3 c), where we have computed the time resolved emission spectrum using the RT-DE scheme. The results differ markedly and show emergent features consistent with non-equilibrium dynamics. For the main renormalized bands, shown in blue dash-dotted lines, we see a picture somewhat similar to the HF-GKBA. Namely there is a renormalization of the band positions compared to equilibrium. Interestingly, for the RT-DE the renormalization effects both the $k\neq 0$ and $k=0$ points very similarly resulting in almost a rigid shift of the bands, compared to the stretching seen in Fig 3 a) and b). This is consistent with renormalized (increased) screening in the photoexcited system which affects the entire band structure and not only the small number states that become occupied by photo-excitation.
@@
In addition, there are two more striking differences between the RT-DE and both TD-HF and the HF-GKBA. The first is a peak appearing just below the renormalized conduction band, consistent with an in-gap electronic state, i.e., an exciton. Secondly we see the appearance of a satellite band appearing just below the position of the equilibrium valence band. Similarly to the finite system in Fig. 2 the inclusion of dynamical correlations through the retarded collision integral allows for the formation of additional spectral features. This is consistent with satellites emerging due to coupling to a bosonic excitation in the system. Here, the satellite is displaced by energy equivalent to the distance between the excitonic feature and the minimum of the renormalized conduction state. The satellites thus can be interpreted as arising due to the simultaneous QP formation and exicton dissociation, i.e., corresponding to an “excitonic replica” of the valence band. The exact interpretation of such features are left for future research, but we note that these are consistent with the exact solutions made for finite systems studied above.
@@
For our tests in finite systems the results for the RT-DE in and out of equilibrium are in excellent agreement with exact results. Not only does the RT-DE correct features picked up by TD-HF and the HF-GKBA but it also captures satellite features that both other methods miss entirely. For our larger system, the RT-DE results provide physically meaningful results. Importantly by including two-body correlations through the self-energy we are able to observe an emergent excitonic peak appearing in the spectral function and energetically well separated from the quasiparticle band.
In conclusion the RT-DE is an efficient scheme for studying the non-equilibrium spectral properties of quantum systems, offering accuracy and a huge reduction in cost relative to the full Kadanoff-Baym equations. Application of this scheme to a more realistic system and comparison of the results to experimental measurements as a validation is underway. Furthermore, deeper investigation into the new spectral features found for periodic system presented here is warranted. Finally, we note that RT-DE can be trivially combined with existing real-time MBPT methods and their extension to more advanced self-energy approximations, such as $GW\Gamma$[43].

### v2

@@ Introduction
In Fig. 3 b) we again show $\mathcal{A}(\omega,t_{p})$, now computed using the HF-GKBA. This result shares features common to TD-HF as we expect because it is dominated by the off-diagonal terms dominated by HF. Firstly, while the HF signatures are diminished, we see a similar stretching of the bands relative to their equilibrium value. One reason for this change comes from an inconsistency between the treatment of the time-diagonal and off-diagonal terms in HF-GKBA, as the HF-GKBA density matrix, $\rho^{\mathrm{HF-GKBA}}(t)$, enters the HF equation for the off-diagonal instead of $\rho^{\mathrm{HF}}(t)$. The density matrix is thus projected on to a set of excited states of the HF Hamiltonian. Such an inconsistency makes it a priori difficult to interpret HF-GKBA spectral results. Importantly the HF-GKBA again lacks additional excitonic features. The picture is thus consistent with TD-HF (albeit with distinct band renormalization ).
@@
Finally we turn to Fig. 3 c), showing the RT-DE results. The results differ markedly and show emergent features. For the main renormalized bands we see a renormalization of the band positions compared to equilibrium. This affects both the $k\neq 0$ and $k=0$ points very similarly resulting in almost a rigid shift of the bands, compared to the stretching seen in Fig 3 a) and b). This is consistent with renormalized (increased) screening in the photoexcited system affecting the entire band structure and not only the small number states that become occupied by photo-excitation.
@@
In addition, there are two more striking differences. The first is a peak appearing just below the renormalized conduction band, consistent with an in-gap electronic state, i.e., an exciton. Secondly we see the appearance of a satellite band appearing just below the position of the equilibrium valence band. Similarly to the finite system in Fig. 2 the inclusion of dynamical correlations through $I^{<}(t,t^{\prime})$ allows for the formation of additional spectral features. This is consistent with satellites emerging due to coupling to a bosonic excitation in the system. Here, the satellite is displaced by energy equivalent to the distance between the excitonic feature and the minimum of the renormalized conduction state. Potentially, one can interpret this as an “excitonic replica”, i.e., simulaneous QP formation and exciton dissociation, similar to other bosonic replicas[42]. The exact interpretation of such features are left for future research, but we note that these are consistent with the exact solutions made for finite systems studied above.
@@
For our tests in finite systems the results for the RT-DE in and out of equilibrium are in excellent agreement with exact results. Not only does the RT-DE correct features picked up by TD-HF and the HF-GKBA but it also captures satellite features that both other methods miss entirely. It is important to highlight that these excellent results are produced using the second Born approximation (representing one of the simplest self-energy forms). Thus, the results and comparisons to exact diagonalization given in this manuscript serve as compelling a proof of principle of the RT-DE. Already our method can capture dynamical correlation effects such as plasmons to a high degree of accuracy and it is natural to expect that extending to more advanced self-energy approximations will only further improve upon the results presented here, an area that is currently being pursued. To show the generality of our approach we have provided a derivation of the RT-DE for the $GW$ self-energy in the SI[26], it’s implementation, testing and comparison is left to a later publication.
For our periodic system, the RT-DE provides physically meaningful results. Importantly by including two-body correlations through $\Sigma(t,t^{\prime})$ we are able to observe an emergent excitonic peak appearing in $\mathcal{A}(\omega,t_{p})$ and energetically well separated from the quasiparticle band.
In conclusion the RT-DE is an efficient scheme for studying the non-equilibrium spectral properties of quantum systems, offering accuracy and a huge reduction in cost relative to the full KBEs. While KBEs and RT-DE will scale similarly in system size, by reducing the cost from cubic to linear in the number of time-steps, we overcome one of the major limitations of the KBEs that has prevented them being widely applied to non-equilibrium first principles simulations. Application of this scheme to a more realistic system and comparison of the results to experimental measurements as a validation is underway. Finally, we note that in principle the RT-DE can be combined with existing real-time MBPT methods and their extension to more advanced self-energy approximations, such as $GW\Gamma$[41]. This extension is currently underway and will open the doors to the accurate and efficient simulation of time-resolved photo-emission experiments.